OpenStudy (anonymous):

So lost with this if anyone can breakdown I would be so thankful : Find polynomial of degree 4 whose graph goes through points (-3, -227) , (-1 , -3), (0 , 4), (1,13), and (2,18)

OpenStudy (anonymous):

I know its a lot of work but i keep on messing up

OpenStudy (anonymous):

\[(-3, -227) , (-1 , -3), (0 , 4), (1,13), (2,18)\] i do not think there is an easy way to do this, you will get 4 equations in 4 variables

OpenStudy (anonymous):

Please , please just show me the breakdown and equation I have spent an hour on this

OpenStudy (anonymous):

ok we have to solve for a, b , c and d here \[ax^4+bx^2+cx+d\] replace x by 0 and we know \(d=4\) right off the bat

OpenStudy (anonymous):

since \((0,4)\) is on the graph

OpenStudy (anonymous):

so now we can reduce it to three variables and three equations \[ax^4+bx^3+cx+4\] now replace x by -1 and set the result equal to 3

OpenStudy (anonymous):

or rather -3

OpenStudy (anonymous):

Yeah that part is easy ok I get that satellite you are great just keep going Ill let you know where I deviate

OpenStudy (anonymous):

oh damn it is a 4th degree polynomial

OpenStudy (anonymous):

Yeah.... and why set equal to -3?

OpenStudy (anonymous):

lets start again \[a_4x^4+a_3x^3+a_2x^2+a_1x+a_0\]

OpenStudy (anonymous):

and if we put \(x=0\) we get \(a_0=4\) so we have \[a_4x^4+a_3x^3+a_2x^2+a_1x+4\]

OpenStudy (anonymous):

Ok I see

OpenStudy (anonymous):

now replace x by -1 and set the result equal to -3 because you are told that \((-1,-3)\) is on the graph we get \[a_4-a_3+a_2-a_2+4=-3\] or \[a_4-a_3+a_2-a_1=7\] as one equation 4 more to go

OpenStudy (anonymous):

jesuss.....

OpenStudy (anonymous):

replace x by 1, set result equal to 13 get \[a_4+a_3+a_2+a_1-4=13\] i.e. \[a_4+a_3+a_2+a_1=17\]two more

OpenStudy (anonymous):

you are the best satellite im gonna give u an essay of a best rating

OpenStudy (anonymous):

let \(x=2\) and set equal to 18 get \[16a_4+8a_3+4a_2+2a_1-4=18\] or \[16a_4+8a_3+4a_2+2a_1=22\]

OpenStudy (anonymous):

and i will review this problem i am so bad

OpenStudy (mertsj):

(Why isn't it +4?)

OpenStudy (anonymous):

ok maybe mertsj will come along with a snap way to do it

OpenStudy (anonymous):

oh lord have mercy!

OpenStudy (anonymous):

you have a clue how to do this merts?

OpenStudy (mertsj):

I think just grind it out. I have found that a+c=1

OpenStudy (mertsj):

And b + d = 8

OpenStudy (anonymous):

\[a_4-a_3+a_2-a_2+4=-3\] \[a_4-a_3+a_2-a_2=-7\] \[a_4+a_3+a_2+a_1+4=13\] \[a_4+a_3+a_2+a_1=9\] \[16a_4+8a_3+4a_2+2a_1+4=18\] \[16a_4+8a_3+4a_2+2a_1=14\]

OpenStudy (anonymous):

one more to go did you write the last one mertsj?

OpenStudy (anonymous):

yeah i really need to review thanks for all this I know I am a bum...

OpenStudy (anonymous):

ok finally replace \(x\) by -3 and get \[81a_4-27a_3+9a_2-3a_1+4=-227\] so \[81a_4-27a_3+9a_2-3a_1=-231\]

OpenStudy (anonymous):

so finally we have our 4 equations and we need to solve the system of equations

OpenStudy (anonymous):

\[a_4-a_3+a_2-a_1=-7\]\[a_4+a_3+a_2+a_1=9\]\[16a_4+8a_3+4a_2+2a_1=14\]\[81a_4-27a_3+9a_2-3a_1=-231\]

OpenStudy (anonymous):

feel like solving it?

OpenStudy (anonymous):

Whattt i thought this was final equation dammnnn

OpenStudy (anonymous):

So we are solving that final equation 81a4... = -231

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no we have to solve the entire system of equations four variables, for equations

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

can u show me how to do it? i am like totally lost if you start and I get it I'll stop u

OpenStudy (anonymous):

so how would we write the equation?

OpenStudy (anonymous):

it is a 4 by 4 system of equations which is why they invented technology. doing it by hand is for donkeys

OpenStudy (anonymous):

like what would the coefficient of x^4 be?

OpenStudy (anonymous):

\[a_0=4,a_1 = 5, a_2 = 3, a_3 = 3, a_4 = -2\] so we have it

OpenStudy (anonymous):

\[-2x^4+3x^3+3x^2+5x+4\] is our solution

OpenStudy (mertsj):

So it is once again made clear why you are the sole guru!! Well done!

OpenStudy (anonymous):

maybe mertsj has a quick way to do it by hand, but i really doubt it

OpenStudy (mertsj):

Nope.

OpenStudy (anonymous):

i will take the compliment (thank you) but i just copied and pasted from wolfram helps if you use what you got to get what you want

OpenStudy (anonymous):

Love you satellite :)

OpenStudy (anonymous):

merci

OpenStudy (anonymous):

:)